Please help ASAP!!!! Homework due at 11pmActivity-Based Supplier Costing Clearsound uses Alpha Electronics and La Paz Company to buy two electronic components used inII. Supplier Data La Paz Company Alpha Electronics 125X 30Y 125X 30Y Unit purchase price $10 $26 $12 $28 Units purchased 15,0
Answer #1

Answer Adpha Slatronds. (25* A Failed tests - 1000 Number of recorder Alpha Electronis = 60 Alpha Electronics Box . Falled teNumber of record = 0 Facted 125x +300+ 125 tov. = 1900 +780+ 10+10 Total = 2600 Number of secoder 3) 125X +300+125 x + 307 ?Recordering componers - 3000 - 300000 100 3000 calculate of unt cost Alpha Electronics La Paz Compar 1254 304 125x__304 1180